# A Comprehensive Overview of Large Language Models

### 20 Feb 2024 v8

@@ III-E LLMs-Powered Agents
AI agents are autonomous entities, capable of planning, decision-making, and performing actions to achieve complex goals. In the older days, AI agents were rule-based, designed for narrow tasks, and had limited capabilities, such as Clippy [216] and Deep Blue [217]. In contrast to this, the LLMs capabilities to respond to dynamic scenarios have made it possible to incorporate them in diverse applications, including LLMs-powered agents [213, 205], where LLMs behave as the brain of agents. LLMs have been incorporated in web agents [155, 156], coding agents [218], tool agents [27, 212], embodied agents [26], and conversational agents [184], without requiring fine-tuning or minimal in some cases. Below we summarize the research in LLMs-based autonomous agents. For a more detailed discussion, please refer to [219, 220].
@@
LLMs Steering Autonomous Agents: LLMs are the cognitive controllers of the autonomous agents. They generate plans, reason about tasks, incorporate memory to complete tasks and adapt the outline depending on the feedback from the environment. Depending on the acquired capabilities of LLMs, many methods fine-tune, propose a better prompting approach, or utilize different modules to enhance agents’ performance. Modules and strategies employed in autonomous agents are briefly discussed below.
@@
Planning and Reasoning: Completing a complex task requires human-like logical thinking, planning necessary steps, and reasoning current and future directions. Prompting methods like chain-of-thoughts [98], tree-of-thoughts [100], and self-consistency [99] are central to agents, eliciting LLMs to reason its actions and choose among different paths for task completion. When LLMs are prompted with task description and sequence of actions, they can accurately generate plan actions without any fine-tuning [221]. Reasoning via planning (RAP) [222] incorporates a re-purposed LLM as a world model to reason about future outcomes and explore alternative paths for task completion. Retroformer [223] uses a retrospective LLM to improve main LLM planning and reasoning capabilities by providing helpful task cues.
@@
Feedback: LLMs in open-loop systems generate plans and assume that the agent will complete it successfully. However, the actual scenario is different with failures and variable responses from the environment. To correctly complete tasks, many methods use LLMs in closed-loop where the action response is provided as feedback to the LLMs to re-assess and update the plan as required [224, 225, 226, 184]. Another direction of research exploits LLMs as reward functions to train reinforcement learning (RL) policies instead of humans [227].
@@
Figure 14: Illustration of parameter-efficient fine-tuning paradigms, where $x$ is input and $h$ is hidden state, figure courtesy [38]. Parallel adapter and LoRA fall in the adapter tuning category
III-F Efficient LLMs
Deploying LLMs in production is expensive. Reducing their running costs while preserving performance is an appealing area of research. This section summarizes the approaches suggested to enhance LLMs efficiency.
III-F1 Parameter Efficient Fine-Tuning
Fine-tuning LLMs with tens or hundreds of billions of parameters, such as GPT-3 (175B), BLOOM (176B), MT-NLG (540B), etc., is computationally-intensive and time-consuming. To avoid complete model fine-tuning, numerous parameter-efficient fine-tuning (PEFT) techniques [40, 236, 41, 38, 39] try to achieve acceptable model fine-tuning performance at reduced costs. As compared to full fine-tuning [237], PEFT performs better in low-resource setups, achieves comparable performance on medium-resource scenarios, and performs worse than full fine-tuning under high-resource availability. An overview of different PEFT approaches is shown in Figure 14.
Adapter Tuning: Adds a few trainable parameters within the transformer block. The adapter layer is a sequence of feature downscaling, non-linearity, and upscaling [101]. Variants of adapter tuning inject adapter layers sequentially [101] and in parallel [38], whereas the mixture of adapter (AdaMix) [238] employs multiple adapter modules in a single layer. AdaMix routes input instances randomly to one of the multiple downscale and upscale modules. The mixture of adapters is averaged out for inference to avoid additional latency. Low-Rank Adaptation (LoRA) [239] learns low-rank decomposed matrices to freeze original weights. Learned weights are fused with the original weights for inference, avoiding latency.
@@ III-F1 Parameter Efficient Fine-Tuning
Bias Tuning: Fine-tuning only bias terms in small to medium training data has been found effective in BitFit [243]. This method achieves full fine-tuning performance for tasks with less training data and comparable performance with larger training data.
III-F2 Quantization
@@
LLMs require extensive computing and memory for inference. Deploying 175B parameter GPT-3 model needs at least 5x80GB A100 GPUs and 350GB of memory to store in FP16 format [44]. Such demanding requirements for deploying LLMs make it harder for smaller organizations to utilize them. Model compression is an effective solution but comes at the cost of degrading performance, especially at large scales greater than 6B. These models exhibit very large magnitude outliers that do not exist in smaller models [244], making it challenging and requiring specialized methods for quantizing LLMs [44, 245].
@@
Post-Training Quantization: Minimal or no training is required in this type of quantization, without significantly compromising the model performance. LLM-8-bit [244] uses full-precision matrix multiplication for weights associated with outlier features and 8-bit for remaining. The lower precision multiplication outputs are converted to FP-16 and concatenated with others. The quantized models have homogenous word embeddings, which may degrade their performance. To fix this, token-level knowledge distillation is employed in [45] along with independent quantization scaling factors for each module due to varying weight distribution. Feature distributions are asymmetric and appear in different channels; outlier suppression [246] shifts and scales per-channel activation distributions for effective quantization. SmoothQuant [44] quantizes activations and weights to INT8 format by smoothing activations and migrating the quantization difficulty toward weights. It multiplies the inverse of the smoothing factor with weights, which introduces a few outliers in the weights but is easier to quantify than unsmoothed activations. OPTQ [245] uses the optimal brain compression (OBC) [247] algorithm to quantize the model layer-by-layer and update weights to compensate for quantization error. To improve speed and performance, OPTQ updates weights in arbitrary order, employs lazy updates, and uses better Cholesky kernels. Outlier-aware weight quantization (OWQ) [248] uses the OPTQ algorithm for quantization but assigns higher precision to vulnerable weights, causing outliers and lower precision for others.
@@
Quantization-Aware Training: To compensate for performance degradation, a quantized model is fine-tuned in quantization-aware training (QAT) [249, 250, 251]. Alphatuning quantizes the model using binary coding quantization (BCQ) [252] and fine-tunes only quantization scaling factors. This approach improves performance over parameter-efficient fine-tuning of a pre-trained model and then quantizing and fine-tuning. Similarly, parameter-efficient and quantization-aware adaptation (PEQA) [253] reduces the precision of fully-connected layers and fine-tunes only quantization scaling parameters. LLM-QAT [251] generates training data from the pre-trained network and trains a quantized student model with knowledge distillation. QLoRA [250] fine-tunes 4-bit quantized pre-trained LLM with LoRA [239] using 4-bit normal float, which shows better performance over 4-bit integer and float.
III-F3 Pruning
@@
Pruning is an alternative approach to quantization to compress model size, thereby reducing LLMs deployment costs significantly. Compared to task-agnostic pruning, task-specific pruning is easily achievable with good performance, where a model is fine-tuned on the downstream task and pruned for faster inference. It is possible to prune LLMs for individual tasks, but the cost of pruning and deploying task-specific models is high. To overcome this, many structured and unstructured pruning methods for LLMs have been proposed to maintain reasonable performance while shrinking the model size [254, 42, 255].
@@
Unstructured Pruning: This kind of pruning removes less important weights without maintaining any structure. Existing LLM pruning methods take advantage of the unique characteristics of LLMs, uncommon for smaller models, where a small subset of hidden states are activated with large magnitude [244]. Pruning by weights and activations (Wanda) [254] prunes weights in every row based on importance, calculated by multiplying the weights with the norm of input. The pruned model does not require fine-tuning, thereby saving computational costs. Outlier weighed layerwise sparsity (OWL) [256] extends Wanda with non-uniform layer pruning. It shows that the number of outliers varies for different layers; therefore, the model should have variable pruning ratios for better performance for every layer. Contrastive pruning (CAP) [43] iteratively prunes the model by training the sparse model using contrastive loss between pre-trained, fine-tuning, and snapshots of previous sparse models to learn task-specific and task-agnostic knowledge.
@@
Structured Pruning: Here, the parameters are removed in groups, rows, columns, or matrices, which speeds up the inference because of effective hardware tensor core utilization [254]. LLM-Pruner [42] employs a 3-stage structured pruning strategy, identifying the groups of hidden states causing each other to activate during forward-pass, keeping important groups and removing less important ones, and fine-tuning the pruned model with LoRA. Sparsity-induced mask learning (SIMPLE) [257] prunes the network using learnable masks. Similarly, another method prunes LLMs by learning masks and removing unimportant rank-1 components of the factorized weight matrix [255].
III-G Multimodal LLMs
@@
Pre-training: This stream of MLLMs intends to support different modalities using unified end-to-end models. For instance, Flamingo [258] applies gated cross-attention to fuse vision and language modalities, which are collected from pre-trained and frozen visual encoder and LLM, respectively. Moreover, BLIP-2 [259] proposes a two-stage strategy to pre-train a Querying Transformer (Q-Former) for the alignment between vision and language modalities: in the first stage, vision-language representation learning is bootstrapped from a frozen visual encoder; and in the second stage, a frozen LLM bootstraps vision-to-language generative learning for zero-shot image-to-text generation. Similarly, MiniGPT-4 [266] also deploys pre-trained and frozen ViT [267], Q-Former and Vicuna LLM [148], while only a linear projection layer needs to be trained for vision and language modalities alignment.
Fine-tuning: Derived from instruction tuning [16] for NLP tasks [20, 16, 92], researchers are now fine-tuning pre-trained LLMs using multimodal instructions. Following this method, LLMs can be easily and effectively extended as multimodal chatbots [266, 260, 29] and multimodal task solvers [268, 30, 269]. The key issue of this stream of MLLMs is to collect multimodal instruction-following data for fine-tuning [58]. To address this issue, the solutions of benchmark adaptation [268, 270, 271], self-instruction [19, 31, 272], and hybrid composition [273, 269] are employed, respectively. To mitigate the gap between the original language modality and additional modalities, the learnable interface is introduced to connect different modalities from frozen pre-trained models. Particularly, the learnable interface is expected to work in a parameter-efficient tuning manner: e.g., LLaMA-Adapter [274] applies an efficient transformer-based adapter module for training, and LaVIN [273] dynamically learns the multimodal feature weights using a mixture-of-modality adapter. Different from the learnable interface, the expert models can directly convert multimodalities into language: e.g., VideoChat-Text [261] incorporates Whisper [275], a speech recognition expert model, to generate the captions of given videos for the understanding of following LLMs.
@@
III-H Summary and Discussion
III-H1 Architecture
@@
Layer Normalization: is found to have a significant effect on the performance and training stability of LLMs. Pre-norm, that is normalizing inputs rather than outputs, is more common among LLMs stabilizing the training [6, 125, 103]. BLOOM [13] and AlexaTM [120] utilize an additional layer normalization before embedding layer to stabilize the training of large-scale models, while the model’s zero-shot generalization ability can be negatively impacted [13]. However, another study [33] finds that pre-norm degrades fine-tuned model performance as compared to post-norm, and there are no stability benefits of pre-norm beyond the 100B scale. Therefore, GLM-130B [33] used deep-norm which is a variant of post-norm for better downstream task performance after fine-tuning.
@@
Positional Encoding: effect performance and training stability of LLMs like other building blocks of a model. BLOOM [13] finds ALiBi outperforming learned and rotary positional encodings. Contrary to this, GLM-130B [33] identifies rotary positional encoding better than ALiBi. So, there is no conclusion in literature about the positional encodings yet.
@@
Parallel Attention: where attention and feed-forward layers are parallel to each other rather than sequential in transformer block has shown to reduce training time by 15%. There is no evidence of performance drop due to this change in literature and used by the models PaLM [15], GPT-NeoX [113], and CodeGen [129].
@@
Multi-Query Attention has shared key and value attention heads in a transformer block while query attention heads are projected as usual. This reduces memory usage and speeds up sampling in autoregressive decoding. No performance degradation has been observed with this change and makes the training efficient allowing larger batch sizes. Multi-query attention is used in [15, 131].
@@
Mixture of Experts: allows easily scaling model to trillion of parameters [128, 116]. Only a few experts are activated during the computation making them compute-efficient. The performance of MoE models is better than the dense models for the same amount of data and requires less computation during fine-tuning to achieve performance similar to the dense models as discussed in [116]. MoE architectures are less prone to catastrophic forgetting, therefore are more suited for continual learning [128]. Extracting smaller sub-models for downstream tasks is possible without losing any performance, making MoE architecture hardware-friendly [128].
@@
Sparse vs Dense Activated: GPT-3 [6] uses sparse transformers [67] whereas GLaM [116] and PanGu-$\sum$ [128] use MoE [117] architecture to lower computational costs and increase the model size and capacity. According to the literature, sparse modules do not degrade the model’s performance [67]. However, more experiments are required to verify this statement.
III-H2 Training Strategies

### 09 Apr 2024 v9

@@ 3.5 LLMs-Powered Agents
AI agents are autonomous entities, capable of planning, decision-making, and performing actions to achieve complex goals. In the early days, AI agents were rule-based, designed for narrow tasks, and had limited capabilities, such as Clippy [217] and Deep Blue [218]. In contrast to this, LLMs abilities to respond to dynamic scenarios have made it possible to incorporate them in diverse applications, including LLMs-powered agents [214, 206], where LLMs behave as the brain of agents. LLMs have been incorporated in web agents [156, 157], coding agents [219], tool agents [27, 213], embodied agents [26], and conversational agents [185], requiring minimal to no fine-tuning". Below we summarize the research in LLMs-based autonomous agents. For a more detailed discussion, please refer to [220, 221].
@@
LLMs Steering Autonomous Agents: LLMs are the cognitive controllers of the autonomous agents. They generate plans, reason about tasks, incorporate memory to complete tasks, and adapt the outline depending on the feedback from the environment. Depending on the acquired capabilities of LLMs, many methods fine-tune, propose a better prompting approach, or utilize different modules to enhance agents’ performance. Modules and strategies employed in autonomous agents are briefly discussed below.
@@
Planning and Reasoning: Completing a complex task requires human-like logical thinking, planning necessary steps, and reasoning current and future directions. Prompting methods like chain-of-thoughts [103], tree-of-thoughts [105], and self-consistency [104] are central to agents, eliciting LLMs to reason its actions and choose among different paths for task completion. When LLMs are prompted with a task description and a sequence of actions, they can accurately generate plan actions without any fine-tuning [222]. Reasoning via planning (RAP) [223] incorporates a re-purposed LLM as a world model to reason about future outcomes and explore alternative paths for task completion. Retroformer [224] uses a retrospective LLM to improve main LLM planning and reasoning capabilities by providing helpful task cues.
@@
Feedback: LLMs in open-loop systems generate plans and assume that the agent will complete them successfully. However, the actual scenario is different with failures and variable responses from the environment. To correctly complete tasks, many methods use LLMs in a closed-loop where the action response is provided as feedback to the LLMs to re-assess and update the plan as required [225, 226, 227, 185]. Another direction of research exploits LLMs as reward functions to train reinforcement learning (RL) policies instead of humans [228].
@@
Figure 14: Illustration of parameter-efficient fine-tuning paradigms, where $x$ is input and $h$ is hidden state, figure courtesy [38]. Parallel adapter and LoRA fall in the adapter tuning category.
3.6 Efficient LLMs
Deploying LLMs in production is expensive. Reducing their running costs while preserving performance is an appealing area of research. This section summarizes the approaches suggested to enhance LLMs’ efficiency.
3.6.1 Parameter Efficient Fine-Tuning
Fine-tuning LLMs with tens or hundreds of billions of parameters, such as GPT-3 (175B), BLOOM (176B), MT-NLG (540B), etc., is computationally intensive and time-consuming. To avoid complete model fine-tuning, numerous parameter-efficient fine-tuning (PEFT) techniques [40, 237, 41, 38, 39] try to achieve acceptable model fine-tuning performance at reduced costs. As compared to full fine-tuning [238], PEFT performs better in low-resource setups, achieves comparable performance on medium-resource scenarios, and performs worse than full fine-tuning under high-resource availability. An overview of different PEFT approaches is shown in Figure 14.
Adapter Tuning: Adds a few trainable parameters within the transformer block. The adapter layer is a sequence of feature downscaling, non-linearity, and upscaling [106]. Variants of adapter tuning inject adapter layers sequentially [106] and in parallel [38], whereas the mixture of adapter (AdaMix) [239] employs multiple adapter modules in a single layer. AdaMix routes input instances randomly to one of the multiple downscale and upscale modules. The mixture of adapters is averaged out for inference to avoid additional latency. Low-Rank Adaptation (LoRA) [240] learns low-rank decomposed matrices to freeze original weights. The learned weights are fused with the original weights for inference, avoiding latency.
@@ 3.6.1 Parameter Efficient Fine-Tuning
Bias Tuning: Fine-tuning only bias terms in small to medium training data has been found effective in BitFit [244]. This method achieves full fine-tuning performance for tasks with less training data and comparable performance with more training data.
3.6.2 Quantization
@@
LLMs require extensive computing and memory for inference. Deploying a 175B parameter GPT-3 model needs at least 5x80GB A100 GPUs and 350GB of memory to store in FP16 format [44]. Such demanding requirements for deploying LLMs make it harder for smaller organizations to utilize them. Model compression is an effective solution but comes at the cost of degraded performance, especially at large scales greater than 6B. These models exhibit very large magnitude outliers that do not exist in smaller models [245], making it challenging and requiring specialized methods for quantizing LLMs [44, 246].
@@
Post-Training Quantization: Minimal or no training is required in this type of quantization, without significantly compromising the model performance. LLM-8-bit [245] uses full-precision matrix multiplication for weights associated with outlier features and 8-bit for remaining features. The lower precision multiplication outputs are converted to FP-16 and concatenated with others. The quantized models have homogenous word embeddings, which may degrade their performance. To fix this, token-level knowledge distillation is employed in [45] along with independent quantization scaling factors for each module due to varying weight distribution. Feature distributions are asymmetric and appear in different channels; outlier suppression [247] shifts and scales per-channel activation distributions for effective quantization. SmoothQuant [44] quantizes activations and weights to INT8 format by smoothing activations and migrating the quantization difficulty toward weights. It multiplies the inverse of the smoothing factor with weights, which introduces a few outliers in the weights but is easier to quantify than unsmoothed activations. OPTQ [246] uses the optimal brain compression (OBC) [248] algorithm to quantize the model layer-by-layer and update weights to compensate for quantization error. To improve speed and performance, OPTQ updates weights in arbitrary order, employs lazy updates, and uses better Cholesky kernels. Outlier-aware weight quantization (OWQ) [249] uses the OPTQ algorithm for quantization but assigns higher precision to vulnerable weights, causing outliers and lower precision for others.
@@
Quantization-Aware Training: To compensate for performance degradation, a quantized model is fine-tuned in quantization-aware training (QAT) [250, 251, 252]. Alpha Tuning quantizes the model using binary coding quantization (BCQ) [253] and fine-tunes only quantization scaling factors. This approach improves performance over parameter-efficient fine-tuning of the pre-trained model. Similarly, parameter-efficient and quantization-aware adaptation (PEQA) [254] reduces the precision of fully-connected layers and fine-tunes only quantization scaling parameters. LLM-QAT [252] generates training data from the pre-trained network and trains a quantized student model with knowledge distillation. QLoRA [251] fine-tunes 4-bit quantized pre-trained LLM with LoRA [240] using a 4-bit normal float, which shows better performance over a 4-bit integer and float.
3.6.3 Pruning
@@
Pruning is an alternative approach to quantization to compress model size, thereby reducing LLMs deployment costs significantly. Compared to task-agnostic pruning, task-specific pruning is easily achievable with good performance, where a model is fine-tuned on the downstream task and pruned for faster inference. It is possible to prune LLMs for individual tasks, but the cost of pruning and deploying task-specific models is high. To overcome this, many structured and unstructured pruning methods for LLMs have been proposed to maintain reasonable performance across all tasks while shrinking the model size [255, 42, 256].
@@
Unstructured Pruning: This kind of pruning removes less important weights without maintaining any structure. Existing LLM pruning methods take advantage of the unique characteristics of LLMs, uncommon for smaller models, where a small subset of hidden states are activated with large magnitude [245]. Pruning by weights and activations (Wanda) [255] prunes weights in every row based on importance, calculated by multiplying the weights with the norm of input. The pruned model does not require fine-tuning, thereby saving computational costs. Outlier weighed layerwise sparsity (OWL) [257] extends Wanda with non-uniform layer pruning. It shows that the number of outliers varies for different layers; therefore, the model should have variable pruning ratios for better performance for every layer. Contrastive pruning (CAP) [43] iteratively prunes the model by training the sparse model using contrastive loss between pre-trained, fine-tuned, and snapshots of previous sparse models to learn task-specific and task-agnostic knowledge.
@@
Structured Pruning: Here, the parameters are removed in groups, rows, columns, or matrices, which speeds up the inference because of effective hardware tensor core utilization [255]. LLM-Pruner [42] employs a 3-stage structured pruning strategy, identifying the groups of hidden states causing each other to activate during the forward-pass, keeping important groups and removing less important ones, and fine-tuning the pruned model with LoRA. Sparsity-induced mask learning (SIMPLE) [258] prunes the network using learnable masks. Similarly, another method prunes LLMs by learning masks and removing unimportant rank-1 components of the factorized weight matrix [256].
3.7 Multimodal LLMs
@@
Pre-training: This stream of MLLMs intends to support different modalities using unified end-to-end models. For instance, Flamingo [259] applies gated cross-attention to fuse vision and language modalities, which are collected from pre-trained and frozen visual encoder and LLM, respectively. Moreover, BLIP-2 [260] proposes a two-stage strategy to pre-train a Querying Transformer (Q-Former) for the alignment between vision and language modalities: in the first stage, vision-language representation learning is bootstrapped from a frozen visual encoder; and in the second stage, a frozen LLM bootstraps vision-to-language generative learning for zero-shot image-to-text generation. Similarly, MiniGPT-4 [267] deploys pre-trained and frozen ViT [268], Q-Former and Vicuna LLM [149], only training the linear projection layer for vision and language modalities alignment.
Fine-tuning: Derived from instruction tuning [16] for NLP tasks [20, 16, 97], researchers are fine-tune pre-trained LLMs using multimodal instructions. Following this method, LLMs can be easily and effectively extended as multimodal chatbots [267, 261, 29] and multimodal task solvers [269, 30, 270]. The key issue of this stream of MLLMs is to collect multimodal instruction-following data for fine-tuning [58]. To address this issue, the solutions of benchmark adaptation [269, 271, 272], self-instruction [19, 31, 273], and hybrid composition [274, 270] are employed, respectively. To mitigate the gap between the original language modality and additional modalities, the learnable interface is introduced to connect different modalities from frozen pre-trained models. Particularly, the learnable interface is expected to work in a parameter-efficient tuning manner: e.g., LLaMA-Adapter [275] applies an efficient transformer-based adapter module for training, and LaVIN [274] dynamically learns the multimodal feature weights using a mixture-of-modality adapter. Different from the learnable interface, the expert models can directly convert multimodalities into language: e.g., VideoChat-Text [262] incorporates Whisper [276], a speech recognition expert model, to generate the captions of given videos for the understanding of following LLMs.
@@
3.8 Summary and Discussion
3.8.1 Architecture
@@
Layer Normalization: The performance and training stability of LLMs are affected significantly by layer normalization. Pre-norm, that is normalizing inputs rather than outputs, is more common among LLMs stabilizing the training [6, 127, 108]. BLOOM [13] and AlexaTM [122] utilize an additional layer normalization before embedding layer to stabilize the training of large-scale models, while the model’s zero-shot generalization ability can be negatively impacted [13]. However, another study [33] finds that pre-norm degrades fine-tuned model performance as compared to post-norm, and there are no stability benefits of pre-norm beyond the 100B scale. Therefore, GLM-130B [33] used deep-norm which is a variant of post-norm for better downstream task performance after fine-tuning.
@@
Positional Encoding: Like other building blocks of the model, positional encoding also affects the performance and training stability of LLMs. BLOOM [13] finds ALiBi outperforms learned and rotary positional encodings. Contrary to this, GLM-130B [33] identifies rotary positional encoding as being better than ALiBi. So, there is no conclusion in the literature about positional encodings yet.
@@
Parallel Attention: In this type of attention, feed-forward and attention layers are parallel to each other rather than sequential in a transformer block. It has been shown to reduce training time by 15%. There is no evidence of performance drop due to this change in the literature and it is used by the models PaLM [15], GPT-NeoX [118], and CodeGen [130].
@@
Multi-Query Attention It has shared key and value attention heads in a transformer block while query attention heads are projected as usual. This reduces memory usage and speeds up sampling in autoregressive decoding. No performance degradation has been observed with this change and it makes the training efficient allowing larger batch sizes. Multi-query attention is used in [15, 132].
@@
Mixture of Experts: This type of architecture enables easily scaling models to trillions of parameters [92, 91]. Only a few experts are activated during the computation making them compute-efficient. The performance of MoE models is better than dense models for the same amount of data and requires less computation during fine-tuning to achieve performance similar to dense models as discussed in [91]. MoE architectures are less prone to catastrophic forgetting, therefore are more suited for continual learning [92]. Extracting smaller sub-models for downstream tasks is possible without losing any performance, making MoE architecture hardware-friendly [92].
@@
Sparse vs Dense Activated: GPT-3 [6] uses sparse transformers [67] whereas GLaM [91] and PanGu-$\sum$ [92] use MoE [121] architectures to lower computational costs and increase the model size and capacity. According to the literature, sparse modules do not degrade the model’s performance [67]. However, more experiments are required to verify this statement.
3.8.2 Training Strategies